# Optimal Transportation by Orthogonal Coupling Dynamics

### v1

@@ Abstract.
Many numerical algorithms and learning tasks rest on solution of the Monge-Kantorovich problem and corresponding Wasserstein distances. While the natural approach is to treat the problem as an infinite-dimensional linear programming, such a methodology severely limits the computational performance due to the polynomial scaling with respect to the sample size along with intensive memory requirements. We propose a novel alternative framework to address the Monge-Kantorovich problem based on a projection type gradient descent scheme. The micro-dynamics is built on the notion of the conditional expectation, where the connection with the opinion dynamics is explored and leveraged to build compact numerical schemes. We demonstrate that the devised dynamics recovers random maps with favourable computational performance. Along with the theoretical insight, the provided dynamics paves the way for innovative approaches to construct numerical schemes for computing optimal transport maps as well as Wasserstein distances.
@@
Emails: mohsen.sadr@psi.ch, P.MohajerinEsfahani@tudelft.nl, Mohammadhossein.Gorji@empa.ch. Mohsen Sadr: Paul Scherrer Institute, Forschungsstrasse 111, CH-5232 Villigen, Switzerland. Peyman Mohajerin Esfehani: The Delft Center for Systems and Control, Delft University of Technology, Delft, The Netherlands. Hossein Gorji: Laboratory for Computational Engineering, Empa, Dübendorf, Switzerland.
@@ 1. Introduction
Background. The importance of the optimal transport lies in its widespread application at different fronts of computational studies, along with its unique theoretical properties. It plays an essential role in machine-learning by offering relevant metrics for comparing different distributions [5, 39, 26]. It has led to a significant progress in the theory of partial differential equations by its by-product, Wasserstein gradient flows [2, 35, 57]. At the same time, existing algorithms to compute such distances and corresponding maps remain expensive and complex. Given the versatile role of optimal transport in different branches of machine-learning [34, 1, 5], analysis [63, 35, 56], density functional theory [18, 24], optimization [43], and inference [32], among others, several numerical approaches have been pursued for its efficient computations. Classical methods were developed based on linear programming [13], gradient descent [19], dynamic flow formulation [10, 62], or elliptic solvers [58]. For efficient computation and improved performance-scaling with respect to the dimension, entropy regularization [25, 51] and its stochastic treatment [7] have been pursued. Approximate approaches using moment formulation [45, 55] of Monge-Kantorovich problem have been proposed, as well.
@@
Despite these progresses, to the best of our knowledge, there is no concise Ordinary Differential Equation (ODE) model, in the sense of particle/agent based models, which leads to the solution of the (un-regularized) Monge-Kantorovich problem. Such a model would be particularly valuable, beyond theoretical and conceptual contributions, for its potential to further enable efficient optimal transport computations. In the opening chapter of Villani’s seminal book [63], a list of coupling dynamics, of ODE or Stochastic Differential Equation (SDE) forms, are reviewed with the intention of creating correlations between two given random variables. While interesting results can be obtained e.g. by the Knoth-Rosenblatt rearrangement [28] or Moser’s coupling [17], such schemes remain limited; e.g. due to coordinate dependency of the Knoth-Rosenblatt rearrangement, or the constraint of small deviation between the marginals for Moser’s coupling.
@@
In principle, finding the optimal transport between two probability spaces is a global optimization in a sense of linear programming on an infinite dimensional space. We are confronted with the question whether there exist local dynamic rules that can lead to the optimal transportation between two sample spaces? An interesting class of dynamic processes is given by orthogonal dynamics. They naturally arise in projecting Hamiltonian systems onto a sub-space, e.g. by the use of conditional expectation [21, 31]. These processes have an interesting feature that the remainder/unresolved portion remains orthogonal to the current state of the resolved variable.
@@
Recent study of Conforti et al. [23] leverages the projection method in the context of Langevin dynamics. Their proposed stochastic dynamics leads to the coupling which converges to the Sinkhorn regularization of the optimal transport problem [25]. Such a stochastic dynamics, which is closely linked to the Schrödinger Bridge [30, 40, 20], gives rise to the Fokker-Planck equation, evolving the joint distribution. It remains to be addressed what happens to the introduced projected Langevin dynamics when the contribution of Brownian motion vanishes, relevant for the (un-regularized) Monge-Kantorovich problem.
@@
The simplest form of the orthogonal dynamics can be constructed by projection of the gradient descent (with respect to the Monge-Kantorovich cost) onto a sub-space where marginals are preserved. In the absence of entropy regularization, this dynamics becomes a good candidate to create optimal transport between two measure spaces. At the level of distribution, one expects such a dynamics gives rise to the Vlasov equation for the evolution of the joint. Despite the theoretical appeal of the coupling through projection, numerical treatment of the conditional expectation can become prohibitive. In general, conditional expectation is solution of the Bayesian regression [46]. On the other hand, interpreting the conditional expectation through distinct data clusters, e.g. in an analogy with opinion dynamics [41], offers non-parametric numerical approach to deal with such coupling dynamics. With these observations in hand and leveraging the recent study [23], we pursue a projected dynamic approach as a concise solution algorithm of the Monge-Kantorovich problem.
Main Contributions. We present the orthogonal coupling dynamics for the Monge-Kantorovich problem. More specifically:
@@
    Main Dynamics. In Sec. 2.1, we introduce the dynamics in Eq. (2.1), motivated by projection of the gradient descent on a marginal-preserving tangent space.
@@
    General Structural Properties. In Sec. 2.2, assuming that the well-posedness of the dynamics is granted, we justify several structural properties for the general cost setting.
@@
    Key Features in $L^{2}$-Monge-Kantorovich. Next, we turn our attention to the common case of $L^{2}$-cost and derive two results in Sec. 2.3.
@@
    Special Examples. In Sec. 3, two special cases of linear maps and elliptic distributions, where explicit results can be obtained, are discussed.
@@
    Numerical Algorithm. In Sec. 4, after motivating the analogy with opinion dynamics, a parameter-free recipe is proposed for computing the conditional expectation.
@@
    Numerical Experiments. We perform several numerical experiments to verify the convergence of the dynamics, as well as its relevance in practical settings, in Sec. 5. We demonstrate that the presented dynamics, equipped with the devised numerical algorithm, successfully recover nonlinear Monge maps. Furthermore, we showcase its performance in distribution learning, data-set classification, and color interpolation.
Our focus remains on basic structural properties of the dynamics along with its computational implications. The aim of the study therefore should be understood as demonstrating the relevance of the orthogonal coupling dynamics, and what it entails, rather than providing a rigorous comprehensive theory, or a thorough numerical analysis, of the presented dynamics itself.
Notation and Setup. Let $\mathcal{P}_{2}(\mathbb{R}^{n})$ be the set of Borel measures over $\mathbb{R}^{n}$ with finite second moment, i.e. $\mathcal{P}_{2}(\mathbb{R}^{n})\mathrel{\mathop{:}}=\{\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})|\int_{\mathbb{R}^{n}}\lVert x\rVert_{2}^{2}\mu(dx)<\infty\}$, where .$\lVert.\rVert_{2}^{2}$ is the usual $L^{2}$ Euclidean norm. Suppose $\Pi(\mu,\nu)$ is the set of all joint measures in $\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ with marginals $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{n})$. We denote by $\pi^{\textrm{opt}}$ the solution of the Monge-Kantorovich problem
[…]
with $c:\mathbb{R}^{2n}\to\mathbb{R}$ as a non-negative smooth function (see e.g. [12] for details). For the standard setting of $c(x,y)=\lVert x-y\rVert_{2}^{2}$, the optimal map $\hat{T}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ exists, if $\mu$ and $\nu$ are absolutely continuous with respect to the Lebesgue measure [2]. It gives $\hat{T}\#\mu=\nu$ (##\## is the push-forward) and accordingly the 2-Wasserstein distance reads
$d^{2}(\mu,\nu)$ $=$ $\int_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\lVert x-y\rVert_{2}^{2}\pi^{\textrm{opt}}_{\mu,\nu}(dx,dy)$ (2)
$=$ $\int_{\mathbb{R}^{n}}\lVert\hat{T}(x)-x\rVert_{2}^{2}\mu(dx)\ .$ (3)
Let $\mathcal{H}=\mathcal{L}^{2}(\Omega,\mathscr{A},P)$ be the Hilbert space comprised of square-integrable $\mathscr{A}$-measurable functions on $\Omega$. Suppose $X_{t},Y_{t}:\Omega\to\mathbb{R}^{n}$ are random processes in $\mathcal{H}$ with the index $t\in[0,+\infty)$ and joint law $p_{t}$. We employ the conditional expectation $\mathbb{E}_{p_{t}}[X_{t}|Y_{t}]$ as the orthogonal projection of $X_{t}$ onto the span of $\sigma(Y_{t})$-measurable functions in $\mathcal{H}$, denoted by $\mathsf{P}_{{p_{t}},{Y_{t}}}[X_{t}]$, where $\sigma(A)$ is the smallest $\sigma$-algebra with respect to which the random variable $A$ is measurable (see e.g. [22, 14] for further details). Accordingly, we have
@@
initialized by $p_{0}\in\Pi(\mu,\nu)$, where the velocity vector $v=(v_{1},v_{2})$ is, in general, a functional of $p_{t}$ and belongs to the tangent space [3]
[…]
and therefore the tangent space is a Hilbert space comprised of gradient vector fields. Besides $\textrm{Tan}_{p_{t}}\mathcal{P}_{2}(\mathbb{R}^{2n})$, we specifically consider two tangent spaces $\textrm{Tan}_{p_{t},1}\Pi(\mu,\nu)$ and $\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$. The former, following [23], reads
$\textrm{Tan}_{p_{t},1}\Pi(\mu,\nu)\mathrel{\mathop{:}}=\left\{v\in\textrm{Tan}_{p_{t}}\mathcal{P}_{2}(\mathbb{R}^{2n}):\int_{\mathbb{R}^{2n}}v\cdot(\nabla h_{1},\nabla h_{2})^{T}p_{t}(dx,dy)=0,\forall h_{1},h_{2}\in C_{c}^{\infty}(\mathbb{R}^{n})\right\}\ ,$ (7)
which gives the set of all velocities that preserve the marginals $\mu$ and $\nu$. Following notation of [23], the latter tangent space is given by
[…]
which is a sub-manifold of $\textrm{Tan}_{p_{t},1}\Pi(\mu,\nu)$, and as shown in Proposition 2.1, it preserves the marginals $\mu$ and $\nu$. Since for $v\in\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$, we have $\mathbb{E}_{p_{t}}[v_{1}(X_{t},Y_{t})|X_{t}]=\mathbb{E}_{p_{t}}[v_{2}(X_{t},Y_{t})|Y_{t}]=0$, the projection operator $\mathsf{S}_{p_{t}}:\textrm{Tan}_{p_{t}}\mathcal{P}_{2}(\mathbb{R}^{2n})\to\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$ is orthogonal to $\mathsf{P}_{p_{t}}$, i.e.

### v2

@@ Abstract.
Many numerical and learning algorithms rely on the solution of the Monge-Kantorovich problem and Wasserstein distances, which provide appropriate distributional metrics. While the natural approach is to treat the problem as an infinite-dimensional linear programming, such a methodology limits the computational performance due to the polynomial scaling with respect to the sample size along with intensive memory requirements. We propose a novel alternative framework to address the Monge-Kantorovich problem based on a projection type gradient descent scheme. The dynamics builds on the notion of the conditional expectation, where the connection with the opinion dynamics is leveraged to devise efficient numerical schemes. We demonstrate that the resulting dynamics recovers random maps with favourable computational performance. Along with the theoretical insight, the proposed dynamics paves the way for innovative approaches to construct numerical schemes for computing optimal transport maps as well as Wasserstein distances.
@@
Emails: msadr@mit.edu, p.mohajerinesfahani@utoronto.ca, Mohammadhossein.Gorji@empa.ch. Mohsen Sadr: Massachusetts Institute of Technology, Cambridge, USA, and Paul Scherrer Institute, Villigen, Switzerland. Peyman Mohajerin Esfahani: University of Toronto, Canada, and Delft University of Technology, the Netherlands. Hossein Gorji: Laboratory for Computational Engineering, Empa, Dübendorf, Switzerland.
@@ 1.1. Background
The importance of the optimal transport lies in its widespread application at different fronts of computational studies, along with its unique theoretical properties. It plays an essential role in machine-learning by offering relevant metrics for comparing different distributions [arjovsky2017wasserstein, kuhn2019wasserstein, dakhmouche2024robust]. It has led to a significant progress in the theory of partial differential equations by its by-product, Wasserstein gradient flows [ambrosio2008gradient, jordan1998variational, santambrogio2017euclidean]. At the same time, existing algorithms to compute such distances and corresponding maps remain expensive and complex. Given the versatile role of optimal transport in different branches of machine-learning [gordaliza2019obtaining, alvarez2018structured, arjovsky2017wasserstein], analysis [villani2009optimal, jordan1998variational, santambrogio2015optimal], density functional theory [buttazzo2012optimal, cotar2013density], optimization [mohajerin2018data], and inference [goldfeld2024statistical], among others, several numerical approaches have been pursued for its efficient computations. Classical methods were developed based on linear programming [bonneel2011displacement], gradient descent [chartrand2009gradient], dynamic flow formulation [benamou2000computational, trigila2016data], or elliptic solvers [saumier2015efficient]. For efficient computation and improved performance-scaling with respect to the dimension, entropy regularization [cuturi2013sinkhorn, peyre2019computational] and its stochastic treatment [ballu2020stochastic] have been pursued. Approximate approaches using moment formulation [mula2024moment, sadr2024wasserstein] of Monge-Kantorovich problem have been proposed, as well.
@@
Despite these progresses, to the best of our knowledge, there is no concise Ordinary Differential Equation (ODE), in the sense of particle/agent based dynamics, which leads to the solution of the (un-regularized) Monge-Kantorovich problem. Such a model would be particularly valuable, beyond theoretical and conceptual contributions, for its potential to further enable efficient optimal transport computations. In the opening chapter of Villani’s seminal book [villani2009optimal], a list of coupling dynamics, of ODE or Stochastic Differential Equation (SDE) forms, are reviewed with the intention of creating correlations between two given random variables. While interesting results can be obtained e.g. by the Knoth-Rosenblatt rearrangement [el2012bayesian] or Moser’s coupling [brenier2003extended], such schemes remain limited; e.g. due to coordinate dependency of the Knoth-Rosenblatt rearrangement, or the constraint of small deviation between the marginals for Moser’s coupling.
@@
In principle, finding the optimal transport map between two probability spaces is a global optimization in the sense of linear programming on an infinite dimensional space. We are confronted with the question whether there exist local dynamic rules that can lead to the optimal transportation between two sample spaces? An interesting class of dynamic processes is given by orthogonal dynamics. They naturally arise in projecting Hamiltonian systems onto a sub-space, e.g. by the use of conditional expectation [chorin2000optimal, givon2005existence]. These processes have an interesting feature that the remainder/unresolved portion remains orthogonal to the current state of the resolved variable.
@@
Recent study of Conforti et al. [conforti2023projected] leverages the projection method in the context of Langevin dynamics. Their proposed stochastic dynamics leads to the coupling which converges to the Sinkhorn regularization of the optimal transport problem [cuturi2013sinkhorn]. Such a stochastic dynamics, which is closely linked to the Schrödinger Bridge [gentil2020dynamical, leonard2013survey, chen2021stochastic], gives rise to the Fokker-Planck equation, evolving the joint distribution. It remains to be addressed what happens to the introduced projected Langevin dynamics when the contribution of Brownian motion vanishes, relevant for the (un-regularized) Monge-Kantorovich problem.
@@
The simplest form of the orthogonal dynamics can be constructed by projection of the gradient descent (with respect to the Monge-Kantorovich cost) onto a sub-space where marginals are preserved. In the absence of entropy regularization, this dynamics becomes a good candidate to create optimal transport between two measure spaces. At the level of distribution, one expects such a dynamics gives rise to the Vlasov equation for the evolution of the joint. Despite the theoretical appeal of the coupling through projection, numerical treatment of the conditional expectation can become prohibitive. In general, conditional expectation is the solution of the Bayesian regression [nadaraya1964estimating]. The Hegselmann-Krause model of opinion dynamics [hegselmann2006truth, li2013consensus] offers non-parametric numerical approach to deal with such coupling dynamics. With these observations in hand and leveraging the recent study [conforti2023projected], we pursue a projected dynamic approach as a concise solution algorithm of the Monge-Kantorovich problem.
1.2. Main Contributions
We propose a numerical scheme for the Monge-Kantorovich problem built on the following steps.
@@ 1.2. Main Contributions
    Coupling Dynamics. We introduce the dynamics in Eq. (14) of section˜2.1, motivated by projection of the gradient descent on a marginal-preserving tangent space.
@@
    General Structural Properties. In section˜2.2, assuming that the well-posedness of the dynamics is granted, we justify several structural properties for the general cost setting.
@@
    Key Features in $L^{2}$-Monge-Kantorovich. Next, we turn our attention to the common case of $L^{2}$-cost and derive two results in section˜2.3.
@@
    Characterization in Gaussian Setting. In section˜4, we study the Riccati dynamics governing the cross-correlation of Gaussian marginals under $L^{2}$-OCD and prove in Proposition 4.1, that the algorithm converges to the true OT if the marginal covariances commute. We further justify that the convergence is exponentially fast in Theorem 4.3.
@@
    Numerical Algorithm. In section˜5, after motivating the analogy with opinion dynamics, a non-parametric recipe is proposed for computing the conditional expectation.
@@
    Numerical Experiments. We perform several numerical experiments to verify the convergence of the dynamics, as well as its relevance in practical settings, in section˜6. We demonstrate that the presented dynamics, equipped with the devised numerical algorithm, successfully recover nonlinear Monge maps. Furthermore, we showcase its performance in distribution learning and color interpolation.
Our focus remains on basic structural properties of the dynamics along with its computational implications. We demonstrate the relevance of the orthogonal coupling dynamics and leave more thorough theoretical and numerical analysis of the dynamics for future studies.
1.3. Notation and Setup
Let $\mathcal{P}_{2}(\Omega)$ be the set of Borel measures over $\Omega$ with a finite second moment, i.e. $\mathcal{P}_{2}(\Omega)\mathrel{\mathop{:}}=\{\mu\in\mathcal{P}(\Omega):\int_{\Omega}\lVert x\rVert_{2}^{2}\mu(dx)<\infty\}$, where $\lVert.\rVert_{2}^{2}$ is the usual $L^{2}$ Euclidean norm, and for $f\in L^{2}(\mu;\mathbb{R}^{n})$ let
$\|f\|_{L^{2}(\mu)}$ $\mathrel{\mathop{:}}=$ $\left(\int_{\mathbb{R}^{n}}\|f(x)\|_{2}^{2}\,\mu(dx)\right)^{1/2}$ (1)
be the $L^{2}$ norm with respect to the measure $\mu$. We write $\mu\otimes\nu$ for the product measure on $\mathbb{R}^{n}\times\mathbb{R}^{n}$. For vectors $A,B\in\mathbb{R}^{n}$, the symbol $A\otimes B$ denotes the rank–one matrix $AB^{\top}$. Suppose $\Pi(\mu,\nu)$ is the set of all joint measures in $\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ with marginals $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{n})$. We denote by $\pi^{\textrm{opt}}$ the solution of the Monge-Kantorovich problem
$\pi^{\textrm{opt}}_{\mu,\nu}$ $\mathrel{\mathop{:}}=$ $\operatorname*{arg\,min}_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{n}\times\mathbb{R}^{n}}c(x,y)\pi(dx,dy)$ (2)
with $c:\mathbb{R}^{2n}\to\mathbb{R}$ as a non-negative smooth function (see e.g. [bogachev2012monge] for details). For the standard setting of $c(x,y)=\frac{1}{2}\lVert x-y\rVert_{2}^{2}$, the optimal map $\hat{T}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ exists, if $\mu$ and $\nu$ are absolutely continuous with respect to the Lebesgue measure [ambrosio2008gradient]. It gives $\hat{T}\#\mu=\nu$ (#\# is the push-forward) and accordingly the 2-Wasserstein distance reads
$d^{2}(\mu,\nu)$ $=$ $\frac{1}{2}\int_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\lVert x-y\rVert_{2}^{2}\pi^{\textrm{opt}}_{\mu,\nu}(dx,dy)$ (3)
$=$ $\frac{1}{2}\int_{\mathbb{R}^{n}}\lVert\hat{T}(x)-x\rVert_{2}^{2}\mu(dx)\ .$ (4)
Let $\mathcal{H}=\mathcal{L}^{2}(\Omega,\mathscr{A},P)$ be the Hilbert space comprised of square-integrable $\mathscr{A}$-measurable functions on $\Omega$. Suppose $X_{t},Y_{t}:\Omega\to\mathbb{R}^{n}$ are random processes in $\mathcal{H}$ with the index $t\in[0,+\infty)$ and joint law $p_{t}$. We employ the conditional expectation $\mathbb{E}_{p_{t}}[X_{t}|Y_{t}]$ as the orthogonal projection of $X_{t}$ onto the span of $\sigma(Y_{t})$-measurable functions in $\mathcal{H}$, denoted by $\mathsf{P}_{{p_{t}},{Y_{t}}}[X_{t}]$, where $\sigma(A)$ is the smallest $\sigma$-algebra with respect to which the random variable $A$ is measurable (see e.g. [chung2000course, borovkov1999probability] for further details). Accordingly, we have
@@ 1.3. Notation and Setup
initialized by $p_{0}\in\Pi(\mu,\nu)$, where the velocity vector $v=(v_{1},v_{2})$ is a functional of $p_{t}$ and belongs to the tangent space [ambrosio2007gradient]
$\textrm{Tan}_{p_{t}}\mathcal{P}_{2}(\mathbb{R}^{2n})$ $\mathrel{\mathop{:}}=$ $\overline{\left\{\nabla h:h\in C_{c}^{\infty}(\mathbb{R}^{2n})\right\}}^{L^{2}(p_{t};\mathbb{R}^{2n})}\ .$ (7)
Note that $\textrm{Tan}_{p_{t}}\mathcal{P}_{2}(\mathbb{R}^{2n})$ is a Hilbert space comprised of gradient vector fields. Besides $\textrm{Tan}_{p_{t}}\mathcal{P}_{2}(\mathbb{R}^{2n})$, we specifically consider two tangent spaces $\textrm{Tan}_{p_{t},1}\Pi(\mu,\nu)$ and $\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$. The former, following [conforti2023projected], reads
$\textrm{Tan}_{p_{t},1}\Pi(\mu,\nu)\mathrel{\mathop{:}}=\left\{v\in\textrm{Tan}_{p_{t}}\mathcal{P}_{2}(\mathbb{R}^{2n}):\int_{\mathbb{R}^{2n}}v\cdot\begin{pmatrix}\nabla h_{1}\\ \nabla h_{2}\end{pmatrix}dp_{t}=0,\forall h_{1,2}\in C_{c}^{\infty}(\mathbb{R}^{n})\right\}\ ,$
which gives the set of all velocities that preserve the marginals $\mu$ and $\nu$. Following notation of [conforti2023projected], the latter tangent space is given by
$\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$ $\mathrel{\mathop{:}}=$ $\left\{v\in\textrm{Tan}_{p_{t}}\mathcal{P}_{2}(\mathbb{R}^{2n}):\begin{pmatrix}\mathbb{E}_{p_{t}}[v_{1}(X,Y)|X]\\ \mathbb{E}_{p_{t}}[v_{2}(X,Y)|Y]\end{pmatrix}=0,\ a.s.\right\},$ (9)
which is a sub-space of $\textrm{Tan}_{p_{t},1}\Pi(\mu,\nu)$. We denote it by $\ker(\mathsf{P}_{p_{t},X})\times\ker(\mathsf{P}_{p_{t},Y})$ for simplicity and as shown in Proposition 2.1, it preserves the marginals $\mu$ and $\nu$. Since for $v\in\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$, we have $\mathbb{E}_{p_{t}}[v_{1}(X_{t},Y_{t})|X_{t}]=\mathbb{E}_{p_{t}}[v_{2}(X_{t},Y_{t})|Y_{t}]=0$, the projection operator $\mathsf{S}_{p_{t}}:\textrm{Tan}_{p_{t}}\mathcal{P}_{2}(\mathbb{R}^{2n})\to\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$ is orthogonal to $\mathsf{P}_{p_{t}}$, i.e.